1.A family desires to have 300,000 available for a daughters college expenses in 20 years. If 8% can be earned on the invested capital, what is the starting amount today?
a. $64,364 
b. $100,916
c. $140,138
d. $150,448


2.One of the famous paintings of Claude Monet was sold in 2016  at auction for $2,000,000,000. If the original selling price in 1926 was the equivalent of $2,500, what is the return on investment (ROI) of this painting?
a. 12.7%.
b. 15.6%.
c. 16.3%. 
d. 18.3%.


3.If $5,000 is invested today into a 4% account and it accumulates to $30,000, what is the number of years required, using logarithms.
a. 31.5 years.
b. 34.7 years.
c. 40.6 years.
d. 45.7 years. 

4.If $20,000 is invested now in an account at Smithtown Mutual Fund, earning 6% annually,  what is the total  amount 25 years later?
a. $81,033.
b. $85,837. 
c. $68,705.
d. $79,599.
Answer #1

Answer:-

(1). Future Value = 300,000

n = 20 years

Rate of Interest = 8%

Present Value = Future Value / (1 + 8%)^20

= 300,000 / (1.08)^20

= 300,000 / 4.661

= 64,363.87 or 64,364 approx.

(2). We need to use CAGR = (Ending value/Beginning value)^(1/no of years) - 1

                                                  = (2000000000 / 2500) ^ (1 / 90) - 1

                                                  = (800,000) ^ (1 / 90) - 1

                                                  = 1.163027 - 1

                                                 = 0.163027

                                                 = 16.30%